study was made of a sample of 25 records of patients seen at a chronic


disease hospital on an


outpatient basis. The mean number of outpatient visits per patient was 4.8,


and the sample standard


deviation was 2. Can it be concluded from these data that the population mean


is greater than four


visits per patient? Let the probability of committing a type I error be .05. What


assumptions are


necessary?

1
Expert's answer
2022-03-20T06:44:27-0400

"H_0:\mu \leqslant 4\\H_1:\mu >4\\T=\sqrt{n}\frac{\bar{x}-\mu}{s}=\sqrt{25}\frac{4.8-4}{2}=2\sim t_{n-1}=t_{24}\\P-value:\\P\left( T>2 \right) =F_{t,24}\left( -2 \right) =0.0285"

Since the P-value is greater than 0.05, the null hypothesis is declined. There is an evidence, that the population mean is greater than four visits per patient.